Harry is 20 years old than Ian. 5 years from now, Ian will be 3/5 as old as Harry. How old is each now?
Answer by jorel555(1290)

Let I be Ian. Then Harry is I+20. So:
I+5=3/5(I+20+5)
5/3(I+5)=I+25
5I+25=3I+75
2I=50
I=25
Ian is 25 right now; Harry is 45. ☺☺☺☺